OpenStudy (peach3113):

@AlexandervonHumboldt2 @TheSmartOne 1. 8(4k-4)=-5k-32 2. -8(-8x-6)=-6x-22 3. 8(1+5x)+5=13+5x 4. -11-5a=6(5a+4) 5. -5(4x-2)=-2(3+6x) 6. 5(2x+6)=-4(-5-2x)+3x

OpenStudy (alexandervonhumboldt2):

you need to simplify them? ok so...............

OpenStudy (alexandervonhumboldt2):

8(4k-4)=-5k-32 distribute: 32k-32=-5k-32

OpenStudy (alexandervonhumboldt2):

37k=0 k=0

OpenStudy (alexandervonhumboldt2):

-8(-8x-6)=-6x-22 64x+48=-6x-22 70x=-70 x=-1

OpenStudy (alexandervonhumboldt2):

8(1+5x)+5=13+5x distribute 8+40x+5=13+5x 0=-35x x=0

OpenStudy (alexandervonhumboldt2):

-11-5a=6(5a+4) -11-5a=30a+24 -35=35a a=-1

OpenStudy (alexandervonhumboldt2):

-5(4x-2)=-2(3+6x) -20x+10=-6-12x -8x=-16 x=2

OpenStudy (alexandervonhumboldt2):

5(2x+6)=-4(-5-2x)+3x 10x+30=20+8x+3x -x=-10 x=10

OpenStudy (peach3113):

Oh gosh thank you Alex.